OpenStudy (anonymous):

if cos0 <0 and sin0= 2/5, find the exact value of tan0

OpenStudy (anonymous):

easiest way to do this is draw a triangle. think of sine as opposite over hypotenuse, so label the opposite side 2 and the hypotenuse 5. all you need is the other leg which you get by pythagoras. \[s^2+2^2=5^2\] \[s^2=5^2-2^2\] \[s=\sqrt{5^2-2^2}=\sqrt{25-4}=\sqrt{21}\]

OpenStudy (anonymous):

yeah i got that

OpenStudy (anonymous):

but how do I get the exact value?

OpenStudy (anonymous):

tan is opposite of adjacent, so you you get \[tan(\theta)=-\frac{5}{\sqrt{21}}\] negative because you are told cosine is negative.

OpenStudy (anonymous):

that is the 'exact value' if you want a decimal approximation you have to use a calculator, but that is not "exact".

OpenStudy (anonymous):

wouldn't i have to multiply by square root of 21 to get rid of the square root?

OpenStudy (anonymous):

if you want to rationalize the denominator that is fine, but rather unnecessary unless you have a math teacher that insists that you do it. you just get \[-\frac{5 \sqrt{21}}{21}\] which doesn't look that much better. in fact it looks worse.

OpenStudy (anonymous):

yes he does insist on it! I wish he didn't. I got everything im just checking all my answers if you don't mind?

OpenStudy (anonymous):

take your time i will check back in a second.

OpenStudy (anonymous):

heres a question im not sure about. Find the length of an arc cut off by a central angle of 1.5 radians in a circle of radius 6 inches.

OpenStudy (anonymous):

radian measure is arc length divided by radius. so if we call the arc length a you have \[\frac{a}{6}=1.5\] \[a=6\times 1.5 = 9\]

OpenStudy (anonymous):

and what is the formula we use for that one?

OpenStudy (anonymous):

and, from a point on the ground, 238 ft from the foot of a vertical tower, the angle of elevation of the top of the tower is 43 degrees. what is the height of the tower to the nearest tenth?

OpenStudy (anonymous):

u there?